### v3

@@ 4 Simulations and Discussion
However, invariant densities for other values of $\lambda$ are not, in general, symmetric. In [22], Hall and Wolff provided numerical estimates for the invariant distributions for the deterministic logistic map for various values of $\lambda$. They demonstrated that any continuous invariant density will have poles (singularities) at the boundary of its support, as well as additional poles at the interior of the support in the case $\lambda<4$. They also show that these invariant distributions were not symmetric. Figure 2 shows numerical approximations to theses deterministic invariant densities for $\lambda=3.87$ and $\lambda=3.95$ These numerical approximations are histograms of 100,000 sampled initial values iterated 20 times. They indicate the singularities in the theoretical densities described in Remark 5 of [22]. These singularities, or poles, correspond to the orbit of the critical point, $f_{\lambda}(1/2)$. In addition, Hall and Wolff observed that the invariant density will be supported on the interval $(\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda)$. This is due to the fact that any point in the interval $(\frac{1}{4}\lambda,1)$ will be mapped into the interval $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}))$, and any point in $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})$ will be mapped to a point in $[\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$. Therefore, after finitely many iterations, any point $x\in(0,1)$ must be mapped into $[\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$, and so the invariant density must be supported on this subinterval. Note that as $\lambda\rightarrow 4$, $\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})\rightarrow 0$ and $\frac{1}{4}\lambda\rightarrow 1$, so the interval of support widens as $\lambda$ approaches 4.

### v4

@@ 4 Simulations and Discussion
However, invariant densities for other values of $\lambda$ are not, in general, symmetric. In [22], Hall and Wolff provided numerical estimates for the invariant distributions for the deterministic logistic map for various values of $\lambda$. They demonstrated that any continuous invariant density will have poles (singularities) at the boundary of its support, as well as additional poles at the interior of the support in the case $\lambda<4$. They also show that these invariant distributions were not symmetric. Figure 2 shows numerical approximations to theses deterministic invariant densities for $\lambda=3.87$ and $\lambda=3.95$ These numerical approximations are histograms of 100,000 sampled initial values iterated 20 times. They indicate the singularities in the theoretical densities described in Remark 5 of [22]. These singularities, or poles, correspond to the orbit of the critical point, $f_{\lambda}(1/2)$. In addition, Hall and Wolff observed that the invariant density will be supported on the interval $(\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda)$. This is due to the fact that any point in the interval $(\frac{1}{4}\lambda,1)$ will be mapped into the interval $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}))$, and any point in $(0,\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})$ will be mapped to a point in $\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$. Therefore, after finitely many iterations, any point $x\in(0,1)$ must be mapped into $\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4}),\frac{1}{4}\lambda]$, and so the invariant density must be supported on this subinterval. Note that as $\lambda\rightarrow 4$, $\frac{1}{4}\lambda^{2}(1-\frac{\lambda}{4})\rightarrow 0$ and $\frac{1}{4}\lambda\rightarrow 1$, so the interval of support widens as $\lambda$ approaches 4.